Grade 11Mechanics

The relation t=√x +3 describes the position of a particle where x is in meters and t is in seconds. The position, when velocity is zero is

Jagjeet Kaur
8 Years agoGrade 11

1 Answer

User
8 Years ago
First of all find v by the relation
For that ..make x as subject in the eq
U wll get x=t2 + 3- 6t
Now for v...
Differentiate with respect to time..
v = dx/dt = 2t + 0 -6= 2t - 6
Now since position at rest is asked so...put v=0
We get..0=2t-6
t = 3
At last...put this t= 3 in the initial.( x = t2 +3 -6t.)
And we get. 3* 3 +2 - 6*3 = -6
Hence x = -6